2. A private research organization studying families in various countries reported the following data for the amount of time 4-year-old children spent alone with their fathers each day.
Country Time with dad (minutes)
Belgium 30
Canada 44
China 54
Finland 50
Germany 36
Nigeria 42
Sweden 46
United States 42

For the above sample, determine the following measures:
a. The mean
b. The standard deviation
c. The mode
1
Expert's answer
2020-08-12T17:07:40-0400

a. The mean is:

"X = (30 + 44 + 54 + 50 + 36 + 42 + 46 + 42)\/8 = 43"minutes.

b. The standard deviation is:

"Sx = (\frac{(30 - 43)^2 + (44 - 43)^2 + (54 - 43)^2 + (50 - 43)^2 + (36 - 43)^2 + (42 - 43)^2 + (46 - 43)^2 + (42 - 43)^2} {8 - 1})^0.5 = 7.56." c. The mode is 42.